1. For a normal curve N (19.73, 1.6), find the area to the right of 18.9. Draw the shaded graph that represents the area of the curve N (19.73, 1.6). Find the corresponding z value and find the area to the right of the z-value. Draw the shaded graph the represents the area under the standard z normal curve.
Answer #1

Solution :

Given that,

mean = = 19.73

standard deviation = =1.6

P(x > 18.9) = 1 - P(x< 18.9)

= 1 - P[(x -) / < (18.9 -19.73) / 1.6]

= 1 - P(z < -0.52)

Using z table

= 1 -  0.3015

=0.6985

area=0.6985